OpenStudy (anonymous):

Determine if the order pair is a solution to the equation. 6x-3y=6

OpenStudy (anonymous):

(2,3)

OpenStudy (anonymous):

I guess 2 is x and 3 is y

OpenStudy (anonymous):

hix do you get this?

OpenStudy (anonymous):

I am trying to use substitution to solve for x and y

OpenStudy (anonymous):

I think there is no solution

OpenStudy (anonymous):

how did you get that?

OpenStudy (radar):

(2,3) is not a solution.

OpenStudy (anonymous):

ok thanx

OpenStudy (radar):

6(2)-3(3)=3 (not 6)

OpenStudy (radar):

Your guess was correct in an ordered pair, the x is first, y second.

OpenStudy (anonymous):

x is always first.?

OpenStudy (anonymous):

right?

OpenStudy (radar):

Yes

OpenStudy (anonymous):

4x-2y=-4 is this a solution? -1=x and 0=y

OpenStudy (radar):

Let me check; 4(-1)-2(0)=-4 Yes!, that is a solution.

OpenStudy (radar):

There are many more solutions, and that was one of them.

OpenStudy (radar):

Just plug in the x value and y value and see if it works.

OpenStudy (anonymous):

5x+y = -15 (2,-5) it equals positive 15 right?

OpenStudy (radar):

5(2)+(-5)=10-5=5 That is not a solution it results in 5, not a -15. Don't forget the y value was a -5 (2,-5)